OpenStudy (anonymous):

A train leaves A at 40 km/hr. At the same time, another train departs from B at a speed of 60 km/h. They reach these respective destinations and turn back immediately towards the starting points. Now if they meet at a distance of 200 km from A, what is the distance between A and B?

OpenStudy (anonymous):

why delete that one

OpenStudy (ranga):



OpenStudy (ranga):

Let x be the distance between A and B. Train A leaves A reaches B, turns around and meets train B at the location marked C. Distance traveled by train A is: x + x - 200 = 2x - 200 Train B leaves B reaches A, turns around and meets train A at the location marked C. Distance traveled by train B is: x + 200. speed = distance / time or time = distance / speed. Time taken by train A = distance / speed = (2x - 200) / 40 Time taken by train B = distance / speed = (x + 200) / 60 But the time taken is the same when the trains meet. (2x - 200) / 40 = (x + 200) / 60 cross multiply and solve for x.